OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3.

OpenStudy (anonymous):

give me equations!! yumyumyum

OpenStudy (goldphenoix):

I forgot.

OpenStudy (anonymous):

i got ALOT @hobbs978

OpenStudy (anonymous):

eliminate all the ones with m not equal to 3

OpenStudy (anonymous):

how ? @hobbs978

OpenStudy (anonymous):

using \[y = mx + b\] put all the equations into that form then see what equations don't have 3 for m

OpenStudy (anonymous):

idkk how @hobbs978

OpenStudy (anonymous):

well here we will take a different approach at this problem then we will solve the the problem with out looking at the provided equation using the equation \[y - y_{1} = m(x-x_{1})\] when you sub in all your given values you have M= 3 y=4 x=2 \[y-4 =3(x-2)\] solving this for y will help you find the equation your looking for

OpenStudy (anonymous):

I honestly don't know how to do that @hobbs978

OpenStudy (goldphenoix):

What do you think you have to do? Give an answer. He'll check if it's right or wrong.

OpenStudy (anonymous):

@GoldPhenoix thank you gold :)

OpenStudy (anonymous):

I'm asking him, not you @GoldPhenoix

OpenStudy (anonymous):

@selena305 Be kind we are all trying to help you here

OpenStudy (anonymous):

Nahh, I dont want its help @hobbs978

OpenStudy (anonymous):

what do you think you should do first if we are trying to solve for y

OpenStudy (anonymous):

subtract ?@hobbs978

OpenStudy (anonymous):

lets take a look at the other side of the equation we see that we have 3(x-2) what do you think we have to do to get () out of our equation